OpenStudy (anonymous):

Cheryl has 16 cupcakes made. She can make 48 cupcakes an hour. A customer asks for 160 cupcakes. How many hours will it take Cheryl to make enough cupcakes for the customer?

OpenStudy (eyust707):

16 made 48 every hour needs 160

OpenStudy (eyust707):

16 + 48 (hours) = 160

OpenStudy (eyust707):

solve for hours

OpenStudy (anonymous):

3 hours?

OpenStudy (eyust707):

yep

OpenStudy (anonymous):

can u help me with sume more

OpenStudy (eyust707):

sure

OpenStudy (anonymous):

Morgan uses 2 oz of dog shampoo to bathe her dog each week. After 4 wk, 34 oz of shampoo remains. What is the total number of weeks that Morgan can use this shampoo before the bottle is empty?

OpenStudy (anonymous):

i think 21

OpenStudy (eyust707):

total amount - 4 * 2 = 34 total amount = 42

OpenStudy (eyust707):

42/2 = 21

OpenStudy (eyust707):

correct

OpenStudy (anonymous):

@eyust707 Very nice explanations!

OpenStudy (eyust707):

thanks! =)

OpenStudy (anonymous):

can u help me with this one

OpenStudy (eyust707):

sure

OpenStudy (eyust707):

lol i love helping people who actually try

OpenStudy (anonymous):

thanks i.g

OpenStudy (eyust707):



OpenStudy (eyust707):

here we want the slope: change in positon/change in time

OpenStudy (anonymous):

330?

OpenStudy (eyust707):

postion went from 495 to 825. thats a change of 825-495 = 330

OpenStudy (eyust707):

yep

OpenStudy (eyust707):

and change in time?

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

2 hours

OpenStudy (eyust707):

no thats how far the train went in two hours

OpenStudy (eyust707):

how far did it go in one hour?

OpenStudy (anonymous):

165

OpenStudy (eyust707):

correct

OpenStudy (eyust707):

330/2 = 165

OpenStudy (anonymous):

so thats my answer

OpenStudy (eyust707):

change in position over change in time equals velocity

OpenStudy (eyust707):

yep

OpenStudy (anonymous):

thanks can u help me with some more?

OpenStudy (eyust707):

sure

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

very close

OpenStudy (eyust707):

notice that they associate the 29 with the y

OpenStudy (eyust707):

and the 6 with the x

OpenStudy (anonymous):

yea

OpenStudy (eyust707):

and also rememeer that slope is: change in y/change in x

OpenStudy (eyust707):

so it will be -7/2 not -2/7

OpenStudy (eyust707):

does that make sense?

OpenStudy (anonymous):

oh yea i was so closed

OpenStudy (eyust707):

yep

OpenStudy (eyust707):

correct all the points on that line always have a y value of -3

OpenStudy (eyust707):

y is always -3

OpenStudy (texaschic101):

HOY -- horizontal line, 0 slope, represented by y = a number

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

VUX -- vertical line, undefined slope, represented by x = a number just in case you wanted to know

OpenStudy (eyust707):

no problem

OpenStudy (eyust707):

correct

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

great job!

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

correct

OpenStudy (eyust707):

haha you got lines down!

OpenStudy (anonymous):

yay i always like double checking my anwers

OpenStudy (eyust707):

thats perfectly fine

OpenStudy (anonymous):

@eyust707

OpenStudy (eyust707):

if he did two deliveries how much money would he earn?

OpenStudy (anonymous):

ummm one sec

OpenStudy (anonymous):

$5

OpenStudy (anonymous):

is it d

OpenStudy (eyust707):

he gets $25 for the day no mater what

OpenStudy (eyust707):

then he gets $2.50 for each delivery

OpenStudy (anonymous):

is it a

OpenStudy (anonymous):

im preety sure its a

OpenStudy (eyust707):

no guessing allowed

OpenStudy (eyust707):

you gotta figure it out

OpenStudy (eyust707):

how much does he get for 2 deliveries

OpenStudy (anonymous):

im not guessing but on a he made 2 deliveries so its $5 + 25 he gets its 30

OpenStudy (eyust707):

25 just for showing up... now ill give you 2.50 for each delivery

OpenStudy (eyust707):

yes!

OpenStudy (anonymous):

so it is a?

OpenStudy (eyust707):

yea

OpenStudy (eyust707):

its A

OpenStudy (anonymous):